OpenStudy (anonymous):

The area of a circular trampoline is 108.94 square feet. What is the radius of the trampoline? Round to the nearest hundredth. A. 10.44 feet B. 5.89 feet C. 34.68 feet D. 3.32 feet

OpenStudy (mathstudent55):

For a circle, A = (pi)r^2 108.94 ft^2 = (pi)r^2 r^2 = (108.94 ft^2)/(pi) r = sqrt(108.94/pi) ft r = 5. 89 ft

OpenStudy (saiyan):

x